# Congruent modular forms and anticyclotomic Iwasawa theory

### v1

@@ 6.1 Definitions
For any field $F$ containing $H_{K}$, there is a decomposition
$H^{1}_{\text{dR}}(A/F)=H^{0}_{\text{dR}}(A/F,\Omega^{1}_{A/F})\oplus H^{1}_{\text{dR}}(A/F,\Omega^{0}_{A/F}).$
Recall our fixed choice of Néron differential $\omega_{A}$, and let $\eta_{A}\in H^{1}_{\text{dR}}(A/F,\Omega^{0}_{A/F})$ such that $\langle\omega_{A},\eta_{A}\rangle=1$ under the algebraic deRham cup product.
@@ 6.2 A $p$-adic Gross-Zagier formula
We recall the following $p$-adic Gross-Zagier formula [5, Theorem 4.9], with the constant term later corrected in the extension to the quaternionic setting due to Magrone [24, Theorem 6.4]. We remark that Theorem 4.9 of CH18 extends the main result of Bertolini-Darmon-Prasanna (see [3, p.1083], [3, Theorem 5.13]) to characters that are ramified at $p$.
@@
Suppose $p=\mathfrak{p}\overline{\mathfrak{p}}$ splits in $K$ and let $f\in S_{2r}(\Gamma_{0}(N))^{\text{new}}$ be a Hecke eigen-newform of weight $2r$. If $\chi=\widehat{\phi}\in\mathfrak{X}_{p^{\infty}}$ is the $p$-adic avatar of an anticyclotomic Hecke character of infinity type $(j,-j)$ with $-r<j<r$ and conductor $p^{n}\mathcal{O}_{K}$ with $n\geq 1$, then
@@
where $\Omega_{p}$ is the $p$-adic period of the canonical elliptic curve $A$ in Section 6.1.
@@
for every anticyclotomic Hecke character $\phi$ of conductor $p^{n}$ and infinity type $(j,-j)$ with $-r<j<r$. Here we used the fact that $v_{p}(\mathfrak{g}(\phi_{\mathfrak{p}}^{-1}))=n/2$ for $n\geq 0$. Under certain hypotheses (see [25, Theorem 5.7], [26, Theorem B]), the $\mu$-invariant $\mu(\mathscr{L}_{\mathfrak{p}}(f))$ vanishes and there is an asymptotic formula:
@@
We now recall the set-up of Section 5. Let $f_{1}\in S_{2r_{1}}(\Gamma_{0}(N_{1}))^{\text{new}},f_{2}\in S_{2r_{2}}(\Gamma_{0}(N_{2}))^{\text{new}}$ be normalized Hecke eigenforms satisfying hypothesis (1). Moreover, suppose that $\chi=\widehat{\phi}\in\mathfrak{X}_{p^{\infty}}$ is the $p$-adic avatar of an anticyclotomic Hecke character of infinity type $(j,-j)$ with $-r<j<r$ and conductor $p^{n}\mathcal{O}_{K}$ with $n\geq 1$. Let $L$ be a number field containing the Hecke eigenvalues of $f_{1}$ and $f_{2}$ as well as the values of $\phi$, and let $\mathcal{W}$ be the ring of integers of the compositum $L_{p}\cdot\widehat{\mathbb{Q}_{p}^{nr}}$. We also denote by $\mathcal{O}_{L_{p}}$ the ring of integers of $L_{p}$. The following Theorem directly follows from Theorem 5.10 and Theorem 6.1.
@@
Suppose that $f_{1},f_{2}$ induce isomorphic semi-simplified $\mod{\varpi}^{m}$ Galois representations: $\bar{\rho}_{f_{1}},\bar{\rho}_{f_{2}}:\operatorname{Gal}(\overline{L}/L)\rightarrow GL_{2}(\mathcal{O}_{L_{p}}/\varpi^{m}\mathcal{O}_{L_{p}})$, where $\varpi$ is the uniformizer of $\mathcal{W}$ and $\mu(\mathscr{L}_{\mathfrak{p}}(f_{1}))=\mu(\mathscr{L}_{\mathfrak{p}}(f_{2}))=0$. Then
$v_{p}(\langle\prod_{\ell\mid N_{1}N_{2}}\mathscr{P}_{\overline{v}}(f_{1})(\chi^{-1})\log_{\mathfrak{p}}(z_{f_{1},\chi})-\prod_{\ell\mid N_{1}N_{2}}\mathscr{P}_{\overline{v}}(f_{2})(\chi^{-1})\log_{\mathfrak{p}}(z_{f_{2},\chi}),\omega_{f}\otimes\omega_{A}^{r-1+j}\eta_{A}^{r-1-j}t^{1-2r}\rangle)$
@@
For a Hecke eigen-newform $f\in S_{2}(\Gamma_{0}(N))^{\text{new}}$ of weight $2$ satisfying Hypothesis (1) and a finite character $\chi$ of conductor $p^{n}$, define
@@
and let $P_{f}(\chi^{-1}):=\pi_{f}(P(\chi^{-1}))$ under the modular parametrization $\pi_{f}:J_{0}(N)\rightarrow A_{f}$. Moroever, let $\omega_{A_{f}}\in H^{0}(A_{f},{\Omega}^{1}_{A_{f}})$ be the differential induced by $f(q)\frac{dq}{q}\in H^{0}(X_{0}(N),{\Omega}^{1}_{X_{0}(N)})$ under the Abel-Jacobi map $\iota:X_{0}(N)\hookrightarrow J_{0}(N)$ and the projection $\pi_{f}$. In the case of weight $2$ forms, we obtain the following extension of [10, Theorem 3.9]:
@@
Suppose that $f_{1}\in S_{2}(\Gamma_{0}(N_{1}))^{\text{new}},f_{2}\in S_{2}(\Gamma_{0}(N_{2}))^{\text{new}}$ induce isomorphic semi-simplified $\mod{\varpi}^{m}$ Galois representations: $\bar{\rho}_{f_{1}},\bar{\rho}_{f_{2}}:\operatorname{Gal}(\overline{L}/L)\rightarrow GL_{2}(\mathcal{O}_{L_{p}}/\varpi^{m}\mathcal{O}_{L_{p}})$, where $\varpi$ is a uniformizer of $L_{p}$. Then
@@ 7 Applications to the Iwasawa Main Conjecture for the BDP Selmer group
where $\mathscr{F}^{+}V$ and $\mathscr{F}^{-}V$ are both $1$-dimensional representations. Let $T$ be a $G_{F}$-stable lattice in $V$ and let $A=V/T$. We also define $\mathscr{F}^{+}T=T\cap\mathscr{F}^{+}V$, $\mathscr{F}^{-}T=T/\mathscr{F}^{+}T$, and $\mathscr{F}^{+}A=\mathscr{F}^{+}V/\mathscr{F}^{+}T$, $\mathscr{F}^{-}A=A/\mathscr{F}^{+}A$.
To define the BDP Selmer group, we recall the following local conditions above $p$, where $M$ is $A,V$ or $T$. Let $F/K$ be a finite extension, and let $v$ be a prime of $F$.
@@
$H^{1}_{\text{Gr}}(F_{v},M):=\begin{cases}\ker\left(H^{1}(F_{v},M)\rightarrow H^{1}(F_{v}^{nr},\mathscr{F}^{-}M)\right)&\text{if }v\mid p,\\ \ker\left(H^{1}(F_{v},M)\rightarrow H^{1}(F_{v}^{nr},M)\right)&\text{if otherwise}.\end{cases}$
@@
$H^{1}_{\mathscr{L}_{v}}(F_{v},M):=\begin{cases}H^{1}(F_{v},M)&\text{if }\mathscr{L}_{v}=\emptyset,\\ H^{1}_{\text{Gr}}(F_{v},M)&\text{if }\mathscr{L}_{v}=\text{Gr},\\ \{0\}&\text{if }\mathscr{L}_{v}=0.\end{cases}$
Let $\Sigma$ be a finite set of primes of $K$ dividing the primes where $V$ is ramified as well as the primes dividing $p\infty$. We will denote by $F_{\Sigma}$ the maximal extension of $F$ unramified outside of the set of primes dividing the primes in $\Sigma$.
@@
For a set of local conditions =$\mathscr{L}=\{\mathscr{L}_{v}\}_{v\mid p}$, we define
$\text{Sel}_{\mathscr{L}}(F,M)=\ker\left(H^{1}(F_{\Sigma}/F,M)\rightarrow\prod_{v\nmid p}\frac{H^{1}(F_{v},M)}{H^{1}_{\text{Gr}}(F_{v},M)}\times\prod_{v\mid p}\frac{H^{1}(F_{v},M)}{H^{1}_{\mathscr{L}_{v}}(F_{v},M)}\right).$

### v2

@@ 6.2 Bloch-Kato logarithm map for a $p$-adic Galois representation
For any $p$-adic field $L$ containing $H_{K,\mathfrak{p}}$, there is a decomposition
$H^{1}_{\text{dR}}(A/L)=H^{0}_{\text{dR}}(A/L,\Omega^{1}_{A/L})\oplus H^{1}_{\text{dR}}(A/L,\Omega^{0}_{A/L}).$
for an elliptic curve $A_{/L}$ with CM by $\mathcal{O}_{K}$. Recall our fixed choice of Néron differential $\omega_{A}$, and let $\eta_{A}\in H^{1}_{\text{dR}}(A/L,\Omega^{0}_{A/L})$ such that $\langle\omega_{A},\eta_{A}\rangle=1$ under the algebraic deRham cup product.
@@ 6.3 A $p$-adic Gross-Zagier formula
Recall the following $p$-adic Gross-Zagier formula [3, Theorem 4.9], with the constant term later corrected in the extension to the quaternionic setting due to Magrone [24, Theorem 6.4]. We remark that Theorem 4.9 of [3] extends the main result of Bertolini-Darmon-Prasanna (see [4, p.1083], [4, Theorem 5.13]) to characters that are ramified at $p$.
@@
[3, Theorem 4.9] Suppose $p=\mathfrak{p}\overline{\mathfrak{p}}$ splits in $K$ and let $f\in S_{2r}(\Gamma_{0}(N))^{\text{new}}$ be a Hecke eigen-newform of weight $2r$. If $\chi=\widehat{\phi}\in\mathfrak{X}_{p^{\infty}}$ is the $p$-adic avatar of an anticyclotomic Hecke character of infinity type (j,j)(j,-j) with r<j<r-r<j<r and conductor $p^{n}\mathcal{O}_{K}$ with $n\geq 1$, then
@@
where $\Omega_{p}$ is the $p$-adic period of the canonical elliptic curve AA in Section 6.1.
@@
for every anticyclotomic Hecke character $\phi$ of conductor $p^{n}$ and infinity type (j,j)(j,-j) with r<j<r-r<j<r. Here we used the fact that $v_{p}(\mathfrak{g}(\phi_{\mathfrak{p}}^{-1}))=n/2$ for $n\geq 0$. Under the extra conditions
[…]
the $\mu$-invariant $\mu(\mathscr{L}_{\mathfrak{p}}(f))$ vanishes [25, Theorem 5.7] (see also [26, Theorem B] for the same statement under slightly different hypotheses) and there is an asymptotic formula:
@@
We now recall the set-up of Section 5. Let $f_{1}\in S_{2r_{1}}(\Gamma_{0}(N_{1}))^{\text{new}},f_{2}\in S_{2r_{2}}(\Gamma_{0}(N_{2}))^{\text{new}}$ be normalized Hecke eigenforms satisfying hypothesis (1). Moreover, suppose that $\chi=\widehat{\phi}\in\mathfrak{X}_{p^{\infty}}$ is the $p$-adic avatar of an anticyclotomic Hecke character of infinity type (j,j)(j,-j) with r<j<r-r<j<r and conductor $p^{n}\mathcal{O}_{K}$ with $n\geq 1$. Let $F$ be a finite extension of $\mathbb{Q}_{p}$ containing the Hecke eigenvalues of $f_{1}$ and $f_{2}$ as well as the values of $\phi$, and let $\mathcal{W}$ be the ring of integers of the compositum $F\cdot\widehat{\mathbb{Q}_{p}^{nr}}$. The following Theorem directly follows from Theorem 5.12 and Theorem 6.1.
@@
Suppose that $f_{1},f_{2}$ induce isomorphic semi-simplified $\mod{\varpi}^{m}$ Galois representations: $\bar{\rho}_{f_{1}},\bar{\rho}_{f_{2}}:G_{\mathbb{Q}}\rightarrow\text{GL}_{2}(\mathcal{O}_{F}/\varpi^{m}\mathcal{O}_{F})$, where $\varpi$ is the uniformizer of $\mathcal{W}$ and $\mu(\mathscr{L}_{\mathfrak{p}}(f_{1}))=\mu(\mathscr{L}_{\mathfrak{p}}(f_{2}))=0$. Then
$v_{p}(\langle\prod_{\ell\mid N_{1}N_{2}}\mathscr{P}_{\overline{v}}(f_{1})(\chi^{-1})\log_{\mathfrak{p}}(z_{f_{1},\chi}),\omega_{f}\otimes\omega_{A}^{r-1+j}\eta_{A}^{r-1-j}t^{1-2r}\rangle$
$-\langle\prod_{\ell\mid N_{1}N_{2}}\mathscr{P}_{\overline{v}}(f_{2})(\chi^{-1})\log_{\mathfrak{p}}(z_{f_{2},\chi}),\omega_{f}\otimes\omega_{A}^{r-1+j}\eta_{A}^{r-1-j}t^{1-2r}\rangle)$
@@
Let $X_{0}(N)$ be the modular curve of level $\Gamma_{0}(N)$ and let $J_{0}(N)$ be its Jacobian. For a Hecke eigen-newform $f\in S_{2}(\Gamma_{0}(N))^{\text{new}}$ of weight $2$ satisfying Hypothesis (1) and a finite character $\chi$ of conductor $p^{n}$, define
@@
[…]
In the case of weight $2$ forms, we obtain the following extension of [10, Theorem 3.9]:
@@
Suppose that $f_{1}\in S_{2}(\Gamma_{0}(N_{1}))^{\text{new}},f_{2}\in S_{2}(\Gamma_{0}(N_{2}))^{\text{new}}$ induce isomorphic semi-simplified $\mod{\varpi}^{m}$ Galois representations: $\bar{\rho}_{f_{1}},\bar{\rho}_{f_{2}}:G_{\mathbb{Q}}\rightarrow\text{GL}_{2}(\mathcal{O}_{F}/\varpi^{m}\mathcal{O}_{F})$, where $\varpi$ is a uniformizer of $F$. Then
@@ 7 Applications to the Iwasawa Main Conjecture for the BDP Selmer group
where $\mathscr{F}^{+}V$ and $\mathscr{F}^{-}V$ are both $1$-dimensional representations. Let $T$ be a $G_{\mathbb{Q}}$-stable lattice in $V$ and let $A=V/T$. We also define $\mathscr{F}^{+}T=T\cap\mathscr{F}^{+}V$, $\mathscr{F}^{-}T=T/\mathscr{F}^{+}T$, and $\mathscr{F}^{+}A=\mathscr{F}^{+}V/\mathscr{F}^{+}T$, $\mathscr{F}^{-}A=A/\mathscr{F}^{+}A$.
To define the BDP Selmer group, we recall the following local conditions above $p$, where $M$ is $A,V$ or $T$. Let $L/K$ be a finite extension of number fields, and let $v$ be a prime of $L$.
@@
$H^{1}_{\text{Gr}}(L_{v},M):=\begin{cases}\ker\left(H^{1}(L_{v},M)\rightarrow H^{1}(L_{v}^{nr},\mathscr{F}^{-}M)\right)&\text{if }v\mid p,\\ \ker\left(H^{1}(L_{v},M)\rightarrow H^{1}(L_{v}^{nr},M)\right)&\text{if otherwise}.\end{cases}$
@@
$H^{1}_{\mathscr{L}_{v}}(L_{v},M):=\begin{cases}H^{1}(L_{v},M)&\text{if }\mathscr{L}_{v}=\emptyset,\\ H^{1}_{\text{Gr}}(L_{v},M)&\text{if }\mathscr{L}_{v}=\text{Gr},\\ \{0\}&\text{if }\mathscr{L}_{v}=0.\end{cases}$
Let $\Sigma$ be a finite set of primes of $K$ dividing the primes where $V$ is ramified as well as the primes dividing $p\infty$. We will denote by $L_{\Sigma}$ the maximal extension of $L$ unramified outside of the set of primes dividing the primes in $\Sigma$.
@@
For a set of local conditions $\mathscr{L}=\{\mathscr{L}_{v}\}_{v\mid p}$, we define
$\text{Sel}_{\mathscr{L}}(L,M)=\ker\left(H^{1}(L_{\Sigma}/L,M)\rightarrow\prod_{v\nmid p}\frac{H^{1}(L_{v},M)}{H^{1}_{\text{Gr}}(L_{v},M)}\times\prod_{v\mid p}\frac{H^{1}(L_{v},M)}{H^{1}_{\mathscr{L}_{v}}(L_{v},M)}\right).$